A charged oil droplet of radius r = 1.5•10-6 m is prevented from falling under gravity by the vertical electric field between two large horizontal plates maintained at a separation of 20mm and a potential difference of 100 V. The density of oil is 920kg/m3.

a) Find the charge on the oil droplet (g = 9.81m/s2).

b) If the polarity of the plates is reversed what is the instantaneous acceleration of the droplet?


1
Expert's answer
2021-04-19T17:10:59-0400

Charge

"m= \rho V=\rho\times \frac{4 \pi}{3}r^3= 920\times \frac{4 \pi}{3}(1.5 \times10^{-6})^3"

"qE=mg; E =\frac{V}{d}=\frac{100}{20 \times10^{-3}}=5000"

"q= \frac{mg}{E}= 920 \times\frac{4 \pi}{3} \times\frac{1.5^3 \times 10^{-18} \times 9.8}{5000}"

"q=2.55 \times 10^{-17} C"

Instantaneous acceleration

"F=qE+mg =mg +mg=2mg"

"ma=2mg"

"a=2g"

"a=2 \times9.81 =19.62 m/s^2"